OpenStudy (anonymous):

solve the substitution method 4x+2y=-4 -8x+y=48 answer must be in ordered pair

OpenStudy (anonymous):

simplify 1st eqn to get 2x + y = -2 frm 2nd eqn y = 48 +8x put this in 1st eqn to get 2x + 48 +8x = -4 solve to get x =10x = - 52 therefore x + -5.2

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

I can show you